OpenStudy (anonymous):

How do I solve Y=x^2+ 4x+ 7 and 2x+ Y+ 2=0 ?

OpenStudy (anonymous):

You want to know what x equals and what y equals, is it?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

There are two method to solve them, one is elimination, one is substitution. First I will tell you substitution.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

It is very easy. The first formula has told you that y=x^2+4x+7. Replace the y in 2x+ Y+ 2=0 with x^2+4x+7, then you get 2x+x^2+4x+7+2=0 x^2+9+6x=0 x^2+3^2+2*3x=0 (x+3)^2=0 \[\sqrt{(x+3)^2}=0\] \[\pm(x+3)=0\] If +(x+3)=0, x=0-3, x=-3; if -(x+3)=0, -x-3=0, -x=3, x=-3 too. So x=-3. Now substitute x=-3 into 2x+ Y+ 2=0. 2*(-3)+y+2=0 -6+2=-y -4=-y y=4 The solution of the equation is x=-3 y=4

OpenStudy (anonymous):

Can you help me with this question? http://openstudy.com/study#/updates/527e0ae5e4b022369a7c11ea

OpenStudy (anonymous):

ill try